A body is thrown vertically upward and attains maximum height in 3 seconds.What is the initial velocity?
1
Expert's answer
2020-07-15T09:35:39-0400

"h=\frac{v_0^2}{2g}" ;


"h=v_0t-\frac{gt^2}{2}" ;


So, we get


"\frac{v_0^2}{2g}=v_0t-\frac{gt^2}{2} \to v_0^2-2gtv_0+g^2t^2=0" or "v_0^2-58.86v_0+866.1249=0\to"


"v_0=29.43m\/s"